A man bought lemons at the rate of 5 for Rs.1 and sold them at 4 for Rs. 1. Find his gain percent.
Answer by stanbon(75887)

A man bought lemons at the rate of 5 for Rs.1 and sold them at 4 for Rs. 1. Find his gain percent.
-----
Each lemon bought at 1/5 Rs
Each lemon sold at 1/4 Rs
---------------
Quantity gain:: (1/4)-(1/5) = 1/20 Rs
-----
Percent gain:: (1/20)/(1/5) = 1/4 = 25%
------------
Cheers,
Stan H.
--------------